Operation Summary for March 22, 2010 - March 24, 2010 |
Monday March 22 The day shift began with Operations monitoring the Tevatron (TeV) store 7695 (100.19E30) and the Recycler stash (271.3E10), with the Antiproton Source stacking (19.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and the Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations prepared the NuMI MI-65 area for an access that started at 8:30 AM and lasted all shift; they were switching to back to neutrino mode. MTest finished running for the day at 10:35 AM. Operations terminated store 7695 at 2:02 PM. Operations put the TeV into shot setup, but a beam-on-beam quench during tune-up brought that to a halt. Recovery lasted through the end of the shift. A Pbar Debuncher RF station (DRF1) power supply failed, causing almost an hour of stacking downtime. There were numerous LRF1 trips. Linac personnel worked on the station. During the evening shift, NuMI was off at the beginning of the shift due to a horn control fault. NuMI personnel traced the problem to the polarity switch work that took place during the day shift. Once this was fixed, NuMI resumed taking beam. TeV recovery and turn-on lasted until 6:18 PM, when Operations put the TeV into shot setup. Operations established store 7698 at 7:42 PM with an average initial luminosity (AIL) of 356.69E30. Store 7698 aborted at 10:56 PM due to a quench indication on a quadrupole power supply (QDD1). There was no quench. The investigation on this lasted through the end of the shift. Operations turned off the MI-10 LCW pumps after the antiprotons were transferred to the TeV. The MI-10 power supply differential pressure switches were reconfigured to run at the resulting lower LCW pressure. The MiniBooNE horn tripped off during this work and would not turn back on. During the investigation into this, operators turned the MI-10 LCW pumps on to get pressure readings at the horn power supply; when they turned the pumps back off, the MI ramp tripped on differential LCW pressure and it took 10-15 minutes to clear the trip and reset. Once the MiniBooNE horn power supply was reconfigured to run at the lower LCW pressure, operators resumed sending beam to MiniBooNE. Tuesday March 23 The midnight shift began with operators still investigating the TeV abort, with Operations monitoring the stash (91.8E10), with Pbar stacking (32.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that the MI ramp tripped off a couple of times on a differential LCW pressure fault at MI-10. Operations called in Mechanical Support personnel to readjust the pressure switches again - this work was finished by 1:56 AM. EE Support personnel traced the TeV QDD1 problem to dirty contacts for the quench detection card. Once this was fixed, Operations put the TeV through dry squeezes and wet squeezes. Operations put the TeV into shot setup at 3:11 AM and established store 7703 at 4:12 AM with an AIL of 173.1E30. During the day shift, FESS personnel turned off and bypassed the MI-10 pond water pump for heat exchanger cleaning. Linac personnel tuned LRF1. Two new Meson MTest experiments (T992 and T979) started taking beam at 2:41 PM. The NuMI beam permit tripped off on a kicker fluorinert temperature fault – cleaning the filters took care of the problem. During the evening shift, Operations disabled the MTest beam at 6:30 PM. Operations terminated store 7703 at 9:52 PM and put the TeV into shot setup. Operations established store 7704 at 11:03PM with an AIL of 368.74E30. Wednesday March 24 The midnight shift began with Operations monitoring store 7704 (242.42E30) and the stash (66.4E10), with Pbar stacking (18.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, a TeV RF station (TRF8) tripped off and caused a slight loss in luminosity and a corresponding change in beam positions. Operations established beam to the MTest experiments at 6:20 AM. There was downtime at the end of the shift due to LRF2 reading back zero even though the station was up and running. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/24/2010 email Fermilab |
| Security, Privacy, Legal | |